# Controlling cooperative emission and superradiance in waveguide-coupled quantum dots

### v1

@@ III Discussion
We present a gated architecture combined with an optimised waveguide design that enables precise control over the QD detunings while maintaining high coupling efficiencies to the waveguide optical modes. This allowed us to create coherence between distant quantum emitters. This architecture including a fast, reversible, and easily scalable tuning method is therefore an important step towards realising large-scale condensed-matter based quantum networks.
IV Methods
IV.1 Nanophotonic devices
The nanophotonic structures are created in a 170 nm GaAs membrane. A layer of QDs is grown at the centre of this membrane using molecular beam epitaxy. The devices are patterned using electron beam lithography, and etched using ICP-RIE. The membrane is grown on top of a 1.1 µm thick sacrificial layer of Al0.6Ga0.4As, which is removed using hydrofluoric acid to leave a suspended structure. The top 30 nm of the GaAs membrane is doped with Carbon at a density of $2\times 10^{19}$ cm-3. The bottom 30 nm of the membrane is doped with Silicon at a density of $2\times 10^{18}$ cm-3. The shallow etch to isolate areas of the device is also defined by electron beam lithography and etched using ICP-RIE to a depth of 40 nm.
IV.2 Theoretical model
In this section we outline our theoretical modelling of superradiant emission from two QDs. More details and derivations of individual terms can be found in Sec. S1 of the Supplementary material [35]. We approximate both QDs as two-level systems such that we can write the Hamiltonian describing both QDs with detuning $\Delta$ as
$H=\frac{\hbar\Delta}{2}(\sigma_{1}^{+}\sigma_{1}^{-}-\sigma_{2}^{+}\sigma_{2}^{-}),$ (2)
where $\sigma_{1/2}^{+}=|e\rangle\langle g|$ ($\sigma_{1/2}^{-}=|g\rangle\langle e|$) is the Pauli spin operator acting in the QD1/2 subspace. The mean energy of both QDs can be set to zero without loss of generality. The time evolution of the system density operator can be described by the Lindblad master equation with Lindblad operators $\mathcal{D}[O]\rho=O{\rho}O^{\dagger}-(O^{\dagger}O\rho-\rho O^{\dagger}O)/2$. We take into account different individual decay rates $\gamma_{1/2}$ for QD1 and QD2 and assume identical dephasing rates $\gamma_{d}$. The finite waveguide quality is captured by the quality factor $\beta$, which is equivalent to standard definitions of the $\beta$-factor for a single QD [22] (for a derivation, see supplementary material [35]). Then, the time evolution of the QDs can be written as
$\dot{\rho}=\frac{1}{i\hbar}$ $[H,\rho]+\beta\mathcal{D}[\Sigma]\rho+(1-\beta)\mathcal{D}[\sqrt{\gamma_{1}}\sigma_{1}^{-}]\rho$
$+(1-\beta)\mathcal{D}[\sqrt{\gamma_{2}}\sigma_{2}^{-}]\rho$
$+\mathcal{D}[\sqrt{\gamma_{d}}\sigma_{1}^{+}\sigma_{1}^{-}]+\mathcal{D}[\sqrt{\gamma_{d}}\sigma_{2}^{+}\sigma_{2}^{-}]\rho,$ (3)
where $\Sigma=\sqrt{\gamma_{1}}\sigma_{1}^{-}+\sqrt{\gamma_{2}}\sigma_{2}^{-}$ encapsulates superradiant decay of two emitters with different decay rates and transition frequencies [39]. To describe the spectral wandering present in the sample we perform a Gaussian-weighted average of realisations with different detunings centered around the observed mean detuning as outlined in Sec. S1.3 of the supplementary material [35].
The lifetime data obtained from TCSPC measurements can be described by the intensity $\langle I(t)\rangle\propto\sum_{\textbf{k}}\langle{a^{\dagger}_{\textbf{k}}}a_{\textbf{k}}\rangle$, where $a_{\textbf{k}}^{\dagger}$ ($a_{\textbf{k}}$) is the creation (annihilation) operator for the waveguide mode with wave vector k. Using the Heisenberg equations of motion, waveguide observables can be related to QD degrees of freedom  [38]. Thus, we use
$\langle I(t)\rangle=\langle\Sigma^{\dagger}\Sigma(t)\rangle/\langle\Sigma^{\dagger}\Sigma(0)\rangle.$ (4)
As initial state we choose the state of only QD1 being excited, $|e_{1},g_{2}\rangle$, and include the timing jitter induced by the above-band excitation via a convolution of the time evolution with a Gaussian function.
In the HBT experiments, the two-times intensity correlation Eq. (1) is measured. Similarly to the intensity, it can be expressed as
$g^{(2)}(\tau)=\frac{\langle\Sigma^{\dagger}(0)\Sigma^{\dagger}(\tau)\Sigma(\tau)\Sigma(0)\rangle_{\rho_{SS}}}{\langle\Sigma^{\dagger}(0)\Sigma(0)\rangle_{\rho_{SS}}\langle\Sigma^{\dagger}(\tau)\Sigma(\tau)\rangle_{\rho_{SS}}},$ (5)
where the index $\langle\bullet\rangle_{\rho_{SS}}$ indicates the expectation value with respect to the steady state density operator. To account for the CW driving applied in these experiments we add additional terms $\mathcal{D}[\sqrt{\gamma_{p}}\sigma_{1}^{+}]\rho$ and $\mathcal{D}[\sqrt{\gamma_{p}}\sigma_{2}^{+}]\rho$ to the master equation (3).
Acknowledgements
@@
Funding: D. H, M.S.S and L.R.W acknowledge support from UK EPSRC (Grant No. EP/V026496/1). M. C. is supported by the Return Program of the State of North Rhine-Westphalia. E.M.G. acknowledges support from UK EPSRC (Grant No. EP/T01377X/1) and from the Leverhulme Trust (RPG-2022-335).
@@
Author contributions: D. H. and L. H performed the experiments. J. W., M. C. and E.M.G. developed the theoretical model. L. H. D. H. and N. M. designed the photonic devices. R. D. fabricated the photonic devices. I. F. and A. V. carried out the growth of the wafer. D. H. and J. W. wrote the manuscript, with contributions from all co-authors. M.S.S. and L.R.W supervised the project.

### v2

@@ Abstract
Discussion   In this work we have presented a gated architecture combined with an optimised waveguide design that enables precise, independent control of emitter wavelengths while maintaining high coupling efficiencies to the waveguide optical modes. This control is fast, reversible and easily scalable. This design was used to create and control coherence between two QDs at a distance 70 times larger than the emission wavelength.
We have performed combined HBT and lifetime measurements on two QDs embedded in a waveguide device. Systematically tuning the QDs in and out of resonance, we demonstrated the transition from a superradiant system to fully independent emitters. At low to zero detuning we demonstrate a decrease in lifetime of about 20%. We validate that this change in lifetime is due to a superradiant rate enhancement by confirming the simultaneous presence of an anti-dip in the HBT data, which indicates the presence of inter-emitter coherences. Additionally, we investigate a region of intermediate detunings, where photon correlations show non-classical behaviour even at detunings of several single-emitter linewidths, while no changes were observed in decay rate. Therefore, our experiments reveal that correlations are still present at very large detunings, where no change in the decay rate is observed, or expected.
@@
D. H, M.S.S and L.R.W acknowledge support from UK EPSRC (Grant No. EP/V026496/1). M. C. is supported by the Return Program of the State of North Rhine-Westphalia. E.M.G. acknowledges support from UK EPSRC (Grant No. EP/T01377X/1) and from the Leverhulme Trust (RPG-2022-335).
@@
D. H. and L. H performed the experiments. J. W., M. C. and E.M.G. developed the theoretical model. D. H, L.R.W, S. S and A. F contributed to the conception of the project. L. H., D. H. and N. M. designed the photonic devices. R. D. fabricated the photonic devices. I. F. and A.K.V. carried out the growth of the wafer. D. H. and J. W. wrote the manuscript, with contributions from all co-authors. M.S.S. and L.R.W supervised the project.